The time for half change in a first order decomposition of a substance A is 60 seconds. Calculate the rate constant. How much of A will be left after 180 seconds?


1
Expert's answer
2021-04-01T02:33:12-0400

Order of a reaction = 1 t1/2 = 60 seconds, k = ?

K = 2.303/60

We know that, k = 2.303/ t1/2

K = 2.303/t1/2 = 0.01155 s^-1

[A] = ?

For the first order reaction k = (2.303/60) log (A0/A)

0.9207 = log 100 – log [A]

 log [A] = log 100 – 0.9207 

log [A] = 2 – 0.9207

log[A] = 1.0973

[A] = antilog of (1.0973) 

[A] = 12.5% or 1/8